Baron Corporation has a target capital structure of 65 percent common stock, 10 percent preferred stock, and 25 percent debt. Its cost of equity is 13 percent, the cost of preferred stock is 8 percent, and the pretax cost of debt is 9 percent. The relevant tax rate is 25 percent. a. What is the company’s WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. What is the aftertax cost of debt?
Answer #1

Aftertax cost of debt=9*(1-tax rate)

=9*(1-0.25)=6.75%

WACC=Respective costs*Respective weight

=(0.65*13)+(0.1*8)+(0.25*6.75)

=10.94%(Approx).